OpenStudy (anonymous):

you reduced your workforce by 1/12 leaving you with 242 people. How many people did you start with?

OpenStudy (alexwee123):

x/12 =242 solve for x

OpenStudy (anonymous):

2904

OpenStudy (alexwee123):

yep ;D

OpenStudy (kropot72):

242 is 11 twelths of the original workforce. So the original workforce is found as follows: \[242\times \frac{12}{11}=?\]